OpenStudy (anonymous):

Help on this one too? "Find 'all' of the solutions for the following equation: 2cosx - 1 = 0"

OpenStudy (shamim):

Cos x=1/2 Right?

OpenStudy (anonymous):

sure? I honestly don't know.

OpenStudy (yamyam70):

\[x = \cos^- 1/2\]

OpenStudy (yamyam70):

to check this substitute that value with x \[2 \cos ( \cos^- 1/2) -1 = 0 \]

OpenStudy (yamyam70):

if the final answer is 0 = 0 , then it is a solution

OpenStudy (anonymous):

ok, so if I'm finding all solutions, would the only solution be 0?

OpenStudy (shamim):

2cos x=1 Right?

OpenStudy (yamyam70):

It depends on what your given is equated to in this case it is equated to 0 , so to know if it is a solution substitute values and put them in calculator , if the result is 0 in your calculator and the given is equated to zero. It is a solution This also applies to whatever your given is equated to., for example if it is equated to 1 result should be 1 , and 1 is equal to 1, 1=1 therefore it is a solution

OpenStudy (michele_laino):

@babygabe97 please note that: 

OpenStudy (anonymous):

ok, I see now.

OpenStudy (shamim):

Feel free to ask if u hv any confusion